OpenStudy (anonymous):

the sum of three consecutive odd integers is 339. what are the integers?

OpenStudy (anonymous):

Think of it this way if you knew one integer was odd, adding two to it would be the next odd integer, such as 1+2=3 Try setting up an equation that uses an idea like that to solve for one variable, and then from that, you can find the other two!

OpenStudy (anonymous):

i need help on this one its too hard

OpenStudy (anonymous):

Okay, I'll give you an equation, but try the rest on your own. You know you're looking for 339, and three different numbers, x represents the lowest, x+2 represents the 2nd, and 2+4 represents the third. 339 = x + (x+2) + (x+4) I only put them in parenthesis, so you can easily identify where they're coming from. Solve for x, and then find the other two integers!